OpenStudy (anikate):

Find the sum of the first 20 terms: 1+3+5+7...

OpenStudy (anonymous):

\[S_n = \frac n2(2a_1 + (n - 1)d)\]

OpenStudy (anikate):

is that a formula?

OpenStudy (anonymous):

a = 1 here and d = 3 - 1 = 2 n = 20 here: \[t_{20} = 1 + (20-1)(2) \implies t_{20} = 39\] Now use: \[Sum = \frac{n}{2}(First + Last)\] n = 20 here..

OpenStudy (anonymous):

Yes. That is the formula if you don't have \(a_n\).

OpenStudy (anonymous):

yea, n = no of terms = 20 a = first term = 1 d = common difference = 2

OpenStudy (anonymous):

First is 1 and Last is 39 n = 20 there.. Plug in the values and find the sum..

OpenStudy (anikate):

ok thx guys I will work it out myself and tell you what I got then u can check it for me.

OpenStudy (anikate):

\[s _{20}=400\]

OpenStudy (anonymous):

Yup :)

OpenStudy (anikate):

ok thx

OpenStudy (anonymous):

np :)